OpenStudy (chancemorris123):

The equation of a line is y = 2x + 3. What is the equation of the line that is parallel to the first line and passes through (2, –1)? A. 4x – 2y = –6 B. y = 2x – 5 C. y = 3x + 4 D. 2x + y = –1

OpenStudy (mathstudent55):

What do you know about the slopes of parallel lines?

OpenStudy (chancemorris123):

anything els? cuz i dont know much

OpenStudy (anonymous):

@mathstudent55 I was just about to post the answer to help him. Oh well, adios!

OpenStudy (mathstudent55):

Parallel lines have the same slope.

OpenStudy (chancemorris123):

yea post it

OpenStudy (mathstudent55):

The given line has equation: y = 2x + 3 That means the slope is 2. The parallel line also has slope 2.

OpenStudy (chancemorris123):

so d? cuz that has 2x

OpenStudy (chancemorris123):

i know its not c

OpenStudy (anonymous):

A and B are parallel to the given equation as @mathstudent55 mentioned in the previous post

OpenStudy (mathstudent55):

The parallel line has equation y = 2x + b Now we need to find b. We use the given point for x and y and solve for b. -1 = 2(2) + b -1 = 4 + b -5 = b Since b = -5, the equation of the parallel line is y = 2x - 5

OpenStudy (mathstudent55):

Which choice is y = 2x - 5?

OpenStudy (chancemorris123):

oh so we know its not c so that is out

OpenStudy (chancemorris123):

b its b

OpenStudy (chancemorris123):

i know it

OpenStudy (mathstudent55):

Correct.

OpenStudy (chancemorris123):

your the man now i can finally sleep i was up all day doing hw

OpenStudy (chancemorris123):

u just gave a child the ability to sleep

OpenStudy (chancemorris123):

god bless

OpenStudy (mathstudent55):

thanks & you're welcome